How many grams of magnesium hydroxide are there in 125 m/ of 3.50 N solution?

1
Expert's answer
2022-05-16T03:28:03-0400

Molarity=Normality/Valence factor=3.5 N / 2= 1.75 M

n(Mg(OH)2) = CM * V = 1.75 M * 0.125 L = 0.219 moles

m(Mg(OH)2) = n(Mg(OH)2) * Mr(Mg(OH)2) = 0.219 moles * 58.33 g/mol = 12.77 g